OpenStudy (anonymous):

find the perimeter and area of the rectangle. both numbers are in units of feet. rectangle shape with 3 on the side and 7 on the bottom. please show all work.

OpenStudy (anonymous):

perimeter of rectangle is 2(l+b) =2(7+3) =20 area = l*b =3*7 =21feet ^2

OpenStudy (anonymous):

m i rite.........

OpenStudy (anonymous):

why is the area 21 feet squared?

OpenStudy (anonymous):

coz its length * breadth area = 7 feet * 3 feet so area = 21 feet^2

OpenStudy (anonymous):

ok. thanks.

OpenStudy (anonymous):

welcum......